### v2

# DAL: A Practical Prior-Free Black-Box Framework for Non-Stationary Bandit Environments

Abstract
We introduce a practical, black-box framework termed Detection Augmenting Learning (DAL) for the problem of non-stationary bandits without prior knowledge of the underlying non-stationarity. DAL is modular, accepting any stationary bandit algorithm as input and augmenting it with a change detector. Our approach is applicable to all common parametric and non-parametric bandit variants. Extensive experimentation demonstrates that DAL consistently surpasses current state-of-the-art methods across diverse non-stationary scenarios, including synthetic benchmarks and real-world datasets, underscoring its versatility and scalability. We provide theoretical insights into DAL’s strong empirical performance on piecewise stationary and drift settings, complemented by thorough experimental validation.
1 Introduction
Bandit models are widely applied in engineering domains, including recommendation systems [35, 36], online advertising [11, 48, 49], dynamic pricing [54], and real-time bidding [23]. Introduced through the multi-armed bandit (MAB) problem [42], many variants have since emerged, which fall into parametric bandits (PB) [4, 19, 22] and non-parametric bandits (NPB) [14, 52]. In a bandit problem, an agent selects actions $A_{t}$ from a compact set $\mathcal{A}\subseteq\mathbb{R}^{d}$ over $T$ rounds, receiving rewards
$X_{t}=f_{t}(A_{t})+\varepsilon_{t},$
where $f_{t}:\mathcal{A}\rightarrow\mathbb{R}$ is the reward function and $\varepsilon_{t}$ is zero-mean sub-Gaussian noise. The goal is to minimize the dynamic regret, using a causal policy $\pi_{t}$ based on past interactions:
$R_{T}:=\mathbb{E}_{A_{t}\sim\pi_{t}}\left[\sum_{t=1}^{T}\max_{a\in\mathcal{A}}f_{t}(a)-f_{t}(A_{t})\right].$
In the case of PB, $f_{t}(A_{t})=\mu\left(\langle\theta_{t},A_{t}\rangle\right)$ where $\theta_{t}$ is an underlying (bounded) parameter unknown to the agent, and $\mu:\mathbb{R}\to\mathbb{R}$ is an injective function. This model includes linear bandits (LBs) with $\mu$ as identity, generalized linear bandits (GLBs), and self-concordant bandits (SCBs) where $\mu$ has the self-concordant property and the noise variance depends on the mean reward [44].
Regarding NPB, we consider the kernelized bandit (KB), under which $f_{t}$ belongs to a reproducing kernel Hilbert space (RKHS) $H_{k}$ induced by a continuous positive semi-definite kernel $k:\ \mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ with $k(x,x)\leq 1$ for all $x\in\mathcal{A}$ and $\|f_{t}\|_{H_{k}}\leq B$. An important quantity for KBs is the maximum information gain (MIG). The MIG at time $t$ is defined as $\gamma_{t}:=\max_{D\subset\mathcal{A}:|D|=t}I(y_{D};f_{D})$, where $I(y_{D};f_{D})$ denotes the mutual information between $f_{D}=[f(x)]_{x\in D}$ and $y_{D}=f_{D}+\varepsilon$, with $\varepsilon\sim\mathcal{N}(0,\lambda v^{2}I)$. It quantifies the reduction in uncertainty about $f$ after observing $y_{D}$ at points $D\subset\mathcal{A}$. For a compact subset $\mathcal{A}\subset\mathbb{R}^{d}$, $\gamma_{T}$ is $\mathcal{O}((\log T)^{d+1})$ for the Squared Exponential (SE) kernel, and $\mathcal{O}(T^{\frac{d(d+1)}{2\nu+d(d+1)}}\log T)$ for the Matérn kernel (with smoothness parameter $\nu$) [53].
The lion’s share of the literature on bandits assumes stationarity—i.e., fixed $f_{t}$ or $\theta_{t}$—but this rarely holds in practice due to evolving user behavior, market conditions, etc.[3, 8, 12, 39, 51]. Non-stationarity is often categorized into two types–gradual drifts and abrupt changes. In the drift model, $f_{t}$ evolves slowly under a variation budget constraint [6, 58]. In contrast, piecewise stationary (PS) models assume abrupt shifts at unknown change-points:
$1=:\nu_{0}<\nu_{1}<\dots<\nu_{N_{T}}<\nu_{N_{T}+1}:=T+1,$
with $f_{t}=f_{t^{\prime}}$ for $t,t^{\prime}\in[\nu_{k},\nu_{k+1})$ and $f_{\nu_{k}-1}\neq f_{\nu_{k}}$ across change-points. Here, $N_{T}$ denotes the total number of changes.
The primary objective of this work is to develop a practical algorithm for non-stationary (NS) bandits that operates without prior knowledge of the underlying non-stationarity. In many real-world applications, changes in the environment—such as shifts in user behavior, emerging trends, or competitive interventions—can occur abruptly rather than gradually. These sudden and decisive shifts can severely impact algorithmic performance, whereas slow, continuous drifts often result in less dramatic degradation and can be handled more gracefully by existing methods. This makes the PS setting not only a natural modeling choice for non-stationarity, but also a case of significant practical relevance. Indeed, as we demonstrate in Section 4.3, algorithms designed for the PS regime can often extend their effectiveness to broader drifting scenarios. For this reason, our focus is firmly placed on the PS setting.
NS bandit algorithms are typically either adaptive—adjusting actions continuously—or restarting, choosing to unlearn and kickstart the learning process at certain times. They may also be prior-based (assuming knowledge of the non-stationarity) or prior-free. Discounted and sliding-window algorithms that weigh recent observations more heavily and forget the distant past are examples of prior-based adaptive approaches, e.g., see [24, 30] for NS-MABs, [13, 45, 57] for NS-LBs, [21, 43, 57] for NS-GLBs, [44, 57] for NS-SCBs, and [16, 63] for NS-KBs. The budget-restarting approaches in [6] for NS-MABs, [61] for NS-LBs, [61] for NS-GLBs, and [63] for NS-KBs are examples of prior-based restarting algorithms. Finally, detection-based restarting algorithms come in both prior-based and prior-free forms. For NS-MABs, [9, 38] constitute such prior-based methods, while [5, 7, 27] are examples of prior-free approaches. For NS-LBs and NS-KBs, see [26]. The most closely related work is that of [27], which addresses the PS-MAB setting and introduces techniques that we build upon in establishing our theoretical insights. However, our setting is much more complex, as it extends to possibly infinite action spaces, where the ideas of [27] are not straightforward to apply. Our framework also applies to the NS-MAB setting, since the NS-LB reduces to NS-MAB as a special case [33]. However, we do not focus on this setting due to the existence of prior approaches [7, 27].
Among prior-free methods, black-box approaches are particularly appealing: they equip any stationary bandit algorithm with non-stationarity handling capabilities. MASTER [58] is the only known order-optimal black-box method for general reinforcement learning. However, as shown in [25], MASTER’s detection mechanism fails to trigger in realistic settings, leading to poor practical performance in experiments. Our work addresses this limitation through the proposal of a black-box framework that is able to achieve good practical performance.
Our Contributions
We introduce a practical, prior-free, detection-based restarting algorithm that is tailored for non-stationary bandits—the first of its kind, to our knowledge. This design choice reflects three key practical insights: (i) prior knowledge of non-stationarity is rarely available; (ii) detection-based methods often enjoy lower worst-case complexity compared to adaptive approaches [40]; and (iii) the black-box paradigm results in significant simplification by reducing the design of algorithms for the non-stationary setting to that of designing how to restart stationary algorithms. In particular, we use a change detector that works with samples obtained from forced exploration to trigger restarts, and combine that with a stationary bandit algorithm. Our design is modular, easy to implement, offering flexibility and adaptability. Extensive empirical evaluation on both synthetic and real-world non-stationary benchmarks demonstrates that our approach consistently outperforms existing state-of-the-art approaches, both prior-free and prior-based. Finally, we provide theoretical insights on the performance of our framework. We show that in the PS setting, under mild conditions, our algorithm’s regret bound matches the state-of-art in PS-LBs and PS-GLBs and improves the current best known bounds for PS-SCBs and PS-KBs. Regarding drifting environments, we offer an intuitive explanation for the conditions under which our approach can achieve good performance, which is validated through additional experiments.
2 The DAL Framework
The DAL framework is a black-box characterized by a modular structure of three components: a non-stationarity detector, a forced exploration scheme, and a bandit algorithm. We provide high-level ideas of the structure of our approach and formally present our framework in Algorithm 1.
Non-Stationarity Detector
To detect changes in the environment, DAL monitors shifts in the representative statistics of the environment, using a general-purpose detector $\mathcal{D}$ for changes in sequences of independent random variables. This distinguishes our approach from methods like MASTER [58], which rely on detecting violations of stationary regret guarantees—an approach not suited for broader detection tasks. In bandit problems, the environment is characterized by reward distributions across actions, so effective change detection requires identifying shifts in these distributions–a premise that strongly correlates with good practical performance, according to [25]. The art lies in employing a detection mechanism that resets learning shortly after a change-point $\nu$, while suppressing the probability of false alarm. We adopt a detection method aligned with [7, 27], grounded in the well-established theory of quickest change detection [41, 55, 56, 59]. DAL keeps a sample history for the rewards of actions that belong in a carefully selected finite subset, and aims to detect changes in those actions.
Forced Exploration
In stationary bandit settings, effective algorithms quickly concentrate on (near-)optimal actions, rarely exploring suboptimal ones. In NS environments, however, this behavior may lead to missed changes on rarely sampled suboptimal actions. To detect such changes, forced exploration is essential. When the action space is large or infinite, uniformly exploring all actions becomes infeasible. To handle infinite action spaces, DAL is able to abstract the detection task through a finite covering set, $\mathcal{A}_{\mathrm{e}}=\{a^{(i)}:i\in[N_{\mathrm{e}}]\}\subseteq\mathcal{A}$, where we specifically use $a^{(i)}$ to denote the $i$-th action in $\mathcal{A}_{\mathrm{e}}$. $\mathcal{A}_{\mathrm{e}}$ must be selected in a way that can guarantee that all significant changes in the reward function can be identified. To do this, the designed $\mathcal{A}_{\mathrm{e}}$ must have at least one action $a^{(i)}$ such that $f_{t}(a^{(i)})$ has changed. DAL is forced to play each action in $\mathcal{A}_{\mathrm{e}}$ once in the first $N_{\mathrm{e}}$ rounds, and then follows the stationary bandit algorithm for the next $\lceil N_{\mathrm{e}}/\alpha_{k}\rceil-N_{\mathrm{e}}$ steps. Here, $\alpha_{k}$ controls the exploration frequency, enabling a balance between performance and reliable change detection.
Bandit Algorithm
The core idea of DAL is to augment any bandit algorithm $\mathcal{B}$ with a change detector $\mathcal{D}$. To facilitate accurate detection, DAL periodically forces $\mathcal{B}$ to sample from the predefined exploration (covering) set $\mathcal{A}_{\mathrm{e}}$. If $\mathcal{D}$ detects a change in the reward distribution associated with any action in $\mathcal{A}_{\mathrm{e}}$, DAL resets the bandit algorithm $\mathcal{B}$ entirely. A key advantage of DAL is its ability to translate strong stationary performance into robust performance under NS conditions. Therefore, by selecting a well-performing stationary algorithm, the DAL framework inherently achieves effective adaptation to NS environments. In fact, the only requirement for DAL’s input stationary algorithm is for it to attain optimal stationary regret performance bounds.
Algorithm 1 Detection Augmented Learning (DAL)
Input: bandit $\mathcal{B}$, detector $\mathcal{D}$, covering set $\mathcal{A}_{\mathrm{e}}$, action set $\mathcal{A}$, horizon $T$, frequencies $\{\alpha_{k}\}_{k=1}^{\infty}$
Initialize: detection $\tau\leftarrow 0$, histories $\mathcal{H}_{a,\mathcal{D}}\leftarrow\emptyset$ $\forall\,a\in\mathcal{A}_{\mathrm{e}}$, counter $k\leftarrow 1$
1:  for $t=1,2,\dots,T$ do
2:     if $(t-\tau-1)\mod\lceil N_{\mathrm{e}}/\alpha_{k}\rceil+1=i\leq N_{\mathrm{e}}$ then
3:        Play action $a^{(i)}$ and receive the reward $X_{t}$.
4:        Add sample $X_{t}$ into history $\mathcal{H}_{a^{(i)},\mathcal{D}}$.
5:        if $\mathcal{D}\left(\mathcal{H}_{a^{(i)},\mathcal{D}}\right)=\mathrm{detection}$ then
6:           Reset the bandit algorithm $\mathcal{B}$.
7:           Empty sample histories for all actions in $\mathcal{A}_{\mathrm{e}}$.
8:           $\tau\leftarrow t$, $k\leftarrow k+1$
9:        end if
10:     else
11:        Run the stationary bandit algorithm $\mathcal{B}$.
12:     end if
13:  end for
3 Practical Performance
In this section, our goal is to establish DAL’s practical effectiveness, by providing experiments on piecewise stationarity and drifting non-stationarity for both synthetic and real-world benchmarks. Since no existing method directly targets the PS-PB and PS-NPB settings with order-optimal guarantees, we compare DAL against the most relevant state-of-the-art algorithms adaptable to PS and drifting settings. We first include MASTER [58], the only other black-box algorithm with order-optimal regret bounds. Although MASTER lacks formal guarantees for SCBs, its recent empirical evaluations [57] support combining it with Logistic-UCB-1 [19] for NS-SCBs. For NS-KBs, despite theoretical results indicating MASTER could be combined with GPUCB [14], MASTER does not attain the state-of-the-art regret bound [26], and due to existing practical issues [25], we omit it from our experiments in NS-KBs. We also evaluate ADA-OPKB [26], the state-of-the-art, prior-free, order-optimal method for NS-LBs and NS-KBs. ADA-OPKB requires extensive parameter tuning for strong performance, posing a challenge in a truly prior-free environment. For fair comparison, however, we tune its parameters to optimize its performance in our evaluation. Finally, we benchmark DAL against two leading prior-based discounted methods: WeightUCB [57], with best theoretical and practical performance for drifting PBs and PS-SCBs, and WGP-UCB [17] for its good performance in drifting KBs. All algorithms use recommended parameters from their original works.
Regarding DAL’s input, in all settings we employ the Generalized Likelihood Ratio (GLR) test [7, 27, 29] as the detector $\mathcal{D}$. Specifically, for the NS-LB setting, we combine DAL with LinUCB [1] and the Gaussian version of the GLR test. For the NS-GLB setting, we employ DAL with GLM-UCB [22] and the Gaussian GLR test. For the NS-SCB setting, we employ DAL with OFUGLB [34], along with the Bernoulli GLR test. Finally, for the NS-KB setting, we combine DAL with REDS [47] along with the Gaussian GLR test. To identify the cover set $\mathcal{A}_{\mathrm{e}}$, we use the information provided in Section 4.1. To implement and practically tune the detectors, we follow the information provided in [27], and provide their structure in the Appendix. Finally, for all settings, we set $\alpha_{k}=10^{-2}\sqrt{kN_{\mathrm{e}}\log(T)/T}$.
3.1 Synthetic Experiments
Figure 1: Dynamic regret vs. environment steps for synthetic experiments. First three columns correspond to the geometric change-points and the final one to the drifting case.
3.1.1 Experimental Parameters
Common parameters
In both NS-PB and NS-KB synthetic experiments, certain parameters are chosen to be the same. Specifically, in all synthetic experiments, the action space comprises $100$ unique actions with dimension $d=10$. These actions are sampled independently from $\mathcal{N}(0,I)$. The horizon is fixed to $T=50000$ and we average the results over $15$ independent runs.
Remark 3.1.
In Algorithm 1, allowing the change detector to use samples from the stationary bandit algorithm $\mathcal{B}$ when $|\mathcal{A}|$ is finite, but only allowing forced exploration over actions in $\mathcal{A}_{\mathrm{e}}$ improves empirical performance. This variation does not affect the theoretical properties of the algorithm, and we employ this variation for our experiments.
NS-PBs
The actions are scaled to lie within an $L$-ball and the underlying parameters $\theta_{t}$ belong to an $S$-ball. Specifically, for NS-LBs and NS-GLBs we have that $S=L=1$, while for NS-SCBs, we have that $L=1$ but $S=3$. Every time a $\theta_{t}$ is initialized or changed, its elements are chosen independently and uniformly in $[-1,1]$, and then are scaled to the $S$-ball. For both GLB and SCB, we select $\mu(x):=(1+e^{-x})^{-1}$. The additive noise is sampled according to $\mathcal{N}(0,0.01)$ at each time-step, while for the SCB setting, we sample the random reward according to $\mathrm{Bernoulli}(\mu(\langle\theta_{t},A_{t}\rangle))$ at time $t$. To select $\mathcal{A}_{\mathrm{e}}$ in the NS-PB setting we use Corollary 4.4.
NS-KBs
Actions are scaled in the $\sqrt{d}$-ball. Regarding the choice of kernel for NS-KBs, we employ the SE kernel with parameter $\ell=0.2$. The noise is sampled according to $\mathcal{N}(0,0.01)$ at each time-step. In the NS-KB setting, we follow a procedure similar to [14, 17]. Specifically, every time we initialize or change the reward function, $f_{t}$ is generated from the RKHS obtained by a discretization of $[-1,1]$ into 200 evenly spaced points. The reward function is generated as $f(\cdot)=\sum_{i=1}^{M}\alpha_{i}k(\cdot,x_{i})$ with $\alpha_{i}$ uniformly distributed in $[-1,1]$ and $M=200$. To identify $\mathcal{A}_{\mathrm{e}}$, we use Corollary 4.4.
3.1.2 Experimental Benchmarks
Piecewise Stationarity
To compare the algorithms in the PS setting, we adopt the geometric change-point model proposed in [25], and independently sample the intervals between the change-points according to a geometric distribution with parameter $\rho=T^{-\xi}$, for $\xi\in\{0.4,0.6,0.8\}$. The expected number of change-points then becomes $\rho T=T^{1-\xi}$, which is sublinear in $T$ for chosen values of $\xi$. Our experiments do not have any restriction on the lengths of the intervals between change-points.
Drifting Non-Stationarity
Regarding comparisons in drifting non-stationarity, we adopt the following drift model: in each run, the reward structure changes linearly over $T$ rounds from an initial value to a final value, where the end-points are chosen the same way described in the beginning of the section. Specifically, we have that
[…]
Experimental Results
Per the results in Figure 1, DAL outperforms the current state-of-art methods in every synthetic experiment. The only competitive method at certain scenarios is WeightUCB, which, unlike DAL, is prior-based. Notably, DAL’s efficacy against the PS benchmarks significantly increases with the decrease in the number of change-points. DAL only abandons the action choices of the stationary bandit algorithm and restarts learning when an efficient change detector flags a mean-shift in rewards; hence, it avoids unnecessary restarts, especially when the intervals between the change-points are long enough for such detectors to correctly flag said changes without false alarms. We suspect that the judicious nature of the restarts is behind its observed superior performance, when changes are further apart. Regarding drifting non-stationarity, DAL significantly outperforms all other methods. In fact, it fares better than both WeightUCB and ADA-OPKB, which not only are known to attain the optimal regret bound in the drift setup, but have also been shown to perform well in practice.
3.2 Real-World Experiments
Figure 2: Results for real-world experiments of Section 3.2 averaged over 15 independent runs.
Click Log Benchmark.
We construct an NS bandit benchmark from the Yahoo! R6A click log dataset. [^1] Yahoo! Front Page Today Module User Click Log Dataset: https://webscope.sandbox.yahoo.com. Following prior works [9, 50], we compute smoothed click-through rates (CTRs) via centered rolling averages over 2000 rounds, average CTRs within each subperiod, and suppress fluctuations below $0.005$. To increase difficulty, we combine actions across 5 days, leading to 64 actions, compress the horizon to $50000$, and multiply final CTRs by 10. We model the resulting environment as an NS-SCB problem, reflecting the logistic reward structure typical in such settings [44].
Stock Market Benchmarks.
NS-KBs have been applied to stock market prediction, and we follow the procedure of [17] to simulate two environments: one using their original data (Quandl stocks) and one constructed from NASDAQ-100 stocks retrieved via the yfinance Python package. [^2] Data retrieved from Yahoo Finance using the publicly available yfinance package. Used solely for non-commercial, academic research purposes. In the Yahoo-based dataset, we retain stocks with sufficient history ($T{=}2000$, approx. 5.5 years) and select the $K=50$ most volatile as actions. Daily closing prices define the reward function, and the empirical price covariance matrix is used as the kernel. To increase difficulty, we add Gaussian noise $\mathcal{N}(0,0.01)$ to the reward at each time-step.
Live Traffic Benchmark.
We construct a NS bandit environment based on the Criteo live traffic dataset [18], following the preprocessing approach of [45] but modeling the problem as an NS-GLB rather than an NS-LB. We estimate the underlying parameter $\theta^{*}$ using logistic regression. Unlike [45], in which the authors employ a single change, we introduce shifts in $\theta^{*}$ via a geometric change-point model with parameter $\xi=0.8$ and extend the horizon to $T=50000$.
Sensor Correlation Benchmark.
We use the Bioliq dataset provided by [31], which contains a week of measurements from 20 sensors in a power plant. We process the reward as [31] and construct an NS-SCB environment with $190$ actions. At each time-step, the reward is $1$ if the last 1000 measurements exceed a threshold of $2.04$, and 00 otherwise. Evaluation is based on cumulative reward.
Ad Recommendation Benchmark.
We evaluate on the Zozo environment, a real-world ad recommender system deployed on an e-commerce platform, introduced by [46]. Using the dataset preprocessed by [31], we construct an NS-GLB environment that captures the dynamics of online ad recommendation. Unlike [31], in which the authors limit the setup to 10 actions due to sparsity, we keep all 80 ads as actions. Following their setup, we assign a reward of 1 to any ad that received at least one user click within a one-second window, and 0 to ads with no clicks. Here, evaluation is based on cumulative reward.
Based on the results in Figure 2, DAL consistently outperforms all state-of-the-art baselines across real-world benchmarks, in both dynamic regret and cumulative reward. We attribute this strong performance to the robustness DAL demonstrates in the synthetic settings, which captured a range of challenging NS scenarios. These findings underscore DAL’s practical effectiveness. In what follows, we provide a theoretical explanation for its performance.
4 Theoretical Insights
4.1 On Effective Detection
Effective non-stationarity detection is crucial for strong practical performance. When selecting a non-stationarity detector, it is essential to ensure accuracy, efficiency, and optimal regret growth properties. Any detector aiming to identify distribution shifts inherently requires a certain number of samples, both before and after the change occurs. Ideally, this sample complexity should scale appropriately to avoid negatively impacting the final regret. To this end, it has been demonstrated that the GLR test achieves a pre- and post-change sample complexity that scales as $\mathcal{O}(\log T)$ [28]. Because logarithmic terms are disregarded in dynamic regret analyses, integrating this detection mechanism properly suggests that optimal regret growth can be achieved.
However, in order to select which samples should be fed into the change detector, one needs to properly design the covering set $\mathcal{A}_{\mathrm{e}}$. When designing $\mathcal{A}_{\mathrm{e}}$, one needs to ensure that it contains actions that can capture the changes in the reward function. However, such changes cannot be allowed to be made arbitrarily small, as no change detector may be able to identify them. $\mathcal{A}_{\mathrm{e}}$ should be designed such that whenever an arbitrary change occurs, reward sequences associated with at least one action in $\mathcal{A}_{\mathrm{e}}$ exhibit an appreciable mean-shift. Define
$\Delta_{\mathrm{c}}\coloneqq\inf_{f\neq f^{\prime}}\max_{a\in\mathcal{A}_{\mathrm{e}}}\ \lvert f(a)-f^{\prime}(a)\rvert.$
Then, $\Delta_{\mathrm{c}}$ captures how well the actions in $\mathcal{A}_{\mathrm{e}}$ can discern between candidate reward functions. According to [27], $\Delta_{\mathrm{c}}$ crucially affects the performance of the GLR test, as its pre- and post- change sample complexity grows with $1/\Delta_{\mathrm{c}}^{2}$. The more discernible the changes are, the easier the detection becomes. Since forced exploration incurs regret, $\mathcal{A}_{\mathrm{e}}$ should ideally be chosen to minimize $N_{\mathrm{e}}$ while maximizing $\Delta_{\mathrm{c}}$. However, this cannot be done since the function $f_{t}$ is unknown. Hence, we provide the ways with which one can ensure that $\mathcal{A}_{\mathrm{e}}$ can lead to detecting the changes in $f_{t}$ in the NS-PB and NS-KB settings. Using these choices of $\mathcal{A}_{\mathrm{e}}$, one can guarantee order-optimal regret in certain cases, as shown in the next section. The proofs of the following statements are given in the Appendix.
Proposition 4.1.
Consider the NS-PB setting. Then, $\mathcal{A}_{\mathrm{e}}$ can be chosen as any arbitrary maximal linearly independent subset of $\mathcal{A}$
Proposition 4.2.
Consider the NS-KB setting and assume w.l.o.g. that $\mathcal{A}\subseteq[0,R]^{d}$. Suppose that there exists an $\tilde{a}\in\mathcal{A}$ such that
$\inf_{f\neq f^{\prime}}|f(\tilde{a})-f^{\prime}(\tilde{a})|>2BL_{u}\delta_{T},$
where $BL_{u}$ is the Lipschitz constant of all $f\in H_{k}(\mathcal{A})$. Let $\mathcal{P}_{T}\subset\mathcal{A}$ correspond to the set of the centers of the balls of an arbitrary $\delta_{T}$-cover. Then, the covering set $\mathcal{A}_{\mathrm{e}}$ can be taken as $\mathcal{P}_{T}$, with $\lvert\mathcal{P}_{T}\rvert\leq\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$.
4.2 On Order-Optimality in Piecewise Stationary Environments
In the PS setting, the minimax regret lower bound under bandit feedback is $\tilde{\Omega}(\sqrt{N_{T}T})$ [24], [^3] We use the $\sim$ in $\tilde{\Omega}(\cdot)$ to hide polylogarithmic factors. which applies across all settings considered in this work, differing only in problem-dependent constants. Under certain conditions on the minimum spacing between change-points stated in [27], our algorithm matches this bound with state-of-the-art dependence on these constants. Specifically, the assumption states that $\nu_{k}-\nu_{k-1}$ should be large enough to trigger the change detector and timely restart.
Let $\ell_{\mathcal{D}}$ and $m_{\mathcal{D}}$ be the latency and the pre-change window length of the detector $\mathcal{D}$ [27], respectively. To characterize DAL’s performance under piecewise stationarity, we employ the methodology of [27], incorporating the regret analysis of the stationary bandit algorithm and that of the change detector. Theorem 4.3 easily follows from Theorem 1 from [27] and a proof sketch is provided in the Appendix.
Theorem 4.3.
For the PS setting of Section 1, consider Algorithm 1 using the GLR test with parameters $\delta_{\mathrm{F}}$ and $\delta_{\mathrm{D}}$, a stationary bandit algorithm $\mathcal{B}$ with regret upper bound $R_{\mathcal{B}}$, a covering set $\mathcal{A}_{\mathrm{e}}$ and forced exploration frequencies $(\alpha_{k})_{k=1}^{\infty}$. Assume $\nu_{1}\geq\lceil N_{\mathrm{e}}/\alpha_{k}\rceil m_{\mathcal{D}}$ and $\nu_{k}-\nu_{k-1}\geq\lceil N_{\mathrm{e}}/\alpha_{k-1}\rceil\ell_{\mathcal{D}}+\lceil N_{\mathrm{e}}/\alpha_{k}\rceil m_{\mathcal{D}}$ for $k\in[2,N_{T}]$. If $R_{\mathcal{B}}(T)=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{T})$ with $p,q\geq 0$, $\delta_{\mathrm{F}}=\delta_{\mathrm{D}}=T^{-\gamma}$ for $\gamma>1$, and $\alpha_{k}=\alpha_{0}\sqrt{kN_{\mathrm{e}}\log\left(T\right)/T}$ with $a_{0}>0$, then DAL’s regret satisfies, $R_{T}=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T}+\sqrt{N_{\mathrm{e}}N_{T}T\log T})$.
Using Theorem 4.3 and Propositions 4.1 and 4.2, we now provide the optimal regret DAL attains.
Corollary 4.4.
Assume that the conditions of Theorem 4.3 hold. In the PS-PB setting, selecting $\mathcal{A}_{\mathrm{e}}$ as in Proposition 4.1, and in the PS-KB setting, selecting $\mathcal{A}_{\mathrm{e}}$ as in Proposition 4.2 with $\delta_{T}:=\frac{Rd^{1/2-2p/d}}{2(C\gamma_{T}^{2q})^{1/d}}$ for some $C>0$, guarantees that DAL attains a regret
$R_{T}=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T}).$
If the base stationary algorithm has order-optimal regret, DAL retains optimality in the PS setting. This also holds when $N_{\mathrm{e}}<d$ or $|\mathcal{A}|<d$ in PS-PBs, and when $N_{\mathrm{e}}<\gamma_{T}$ or $|\mathcal{A}|<\gamma_{T}$ in PS-KBs.
Remark 4.5.
The assumption on the change-points is necessary to prove the order-optimality, but it is not for practical performance. None of our experiments enforced this assumption, and DAL dominated in both the synthetic and the real-world simulations as shown in Section 3.
The assumption on the minimum separation between change-points essentially requires scaling as $\tilde{\mathcal{O}}(\sqrt{T/k})$. However, this condition primarily emerges from a conservative proof technique used in [27], where missed detections are aggregated into a single adverse event. Practically, and as corroborated by our experiments in Section 3, this assumption is often violated without negatively impacting the regret performance—even under scenarios with frequent and arbitrarily placed change-points (e.g. $\xi=0.4$). We suspect that this resilience arises because the GLR test, while potentially missing isolated short intervals, reliably detects subsequent changes when stationary segments exceed the threshold length. Even if a change is entirely missed during a segment shorter than $\tilde{\mathcal{O}}(\sqrt{T/k})$, the resulting regret remains under that order. Conversely, when the assumption holds, optimal regret is provably guaranteed. Thus, the required separation threshold acts as a practical ”sweet spot”: segments longer than this threshold are detected reliably, ensuring optimal performance, while shorter segments incur minimal regret, thereby preserving overall optimal regret guarantees.
In line with the black-box design philosophy, Corollary 4.4 enables regret upper bounds across all settings considered, with flexibility in the choice of stationary bandit algorithms. When using specific stationary algorithms from Section 3, DAL matches the state-of-the-art regret bounds in PS-LBs and PS-GLBs at $\tilde{\mathcal{O}}(d\sqrt{N_{T}T})$. More notably, DAL improves the best known bounds in the PS-SCB and PS-KB settings. For PS-SCBs, the strongest, prior-based, bound is due to WeightUCB [57], which achieves $\tilde{\mathcal{O}}(d^{2/3}T^{2/3}N_{T}^{1/3})$ [^4] While MASTER [58] may be extendable to PS-SCBs, no regret bound is currently known.. DAL improves this to $\tilde{\mathcal{O}}(d\sqrt{N_{T}T})$ with our algorithmic choices. Although this matches the bound in [44], their analysis relies on substantially stronger assumptions than those in [27]. For PS-KBs, the prior-free ADA-OPKB [26] achieves $\tilde{\mathcal{O}}(\sqrt{d\gamma_{T}N_{T}T})$, while DAL improves this to $\tilde{\mathcal{O}}(\sqrt{\gamma_{T}N_{T}T})$. This highlights the interesting feature of DAL: the order-wise dependence on problem parameters from the stationary setting seamlessly transfers to the PS setting without degradation. A more detailed comparison of regret bounds is provided in the Appendix.
4.3 On Drifting Environments
Based on the previous section, at first glance, one can expect that DAL is not able to handle drifting non-stationarity. Our results in Section 3 naturally lead us to ask when and why DAL performs well in drifting environments. As a first step to study this, we perform another experiment with LBs. Specifically, the parameter $\theta_{t}$ in each time-step $t$ evolves randomly as follows,
$\theta_{t+1}:=\theta_{t}+\zeta_{t+1}$
where $\zeta_{t+1}\in\mathbb{R}^{d}$ is chosen uniformly over a $\delta$-ball. If the resulting $\theta_{t+1}$ violates the norm-bound $S$, we disregard that choice of $\zeta_{t+1}$ and sample again. We sample $\varepsilon_{t}\sim\mathcal{N}(0,0.1)$ at each $t$. We compare the cumulative dynamic regret up to time $T$ of DAL+LinUCB and WeightUCB over a range of $\delta$’s in Figure 3. The remaining parameters are chosen to be the same as those in Section 3, with the exception of $d=5$. The DAL algorithm performs better than WeightUCB for smaller values of $\delta$, but the conclusion reverses upon increasing $\delta$.
Figure 3: Final dynamic regret vs. radius of change $\delta$: Drifting LBs.
We now shed light into our hypothesis behind the observations from Figure 3. For playing an action $a\in\mathcal{A}$ at time $t+1$, we get the random reward,
[…]
If the governing parameter does not change, then the corresponding reward from playing action $a$ would have been $X^{\prime}_{t+1}=\langle\theta_{t},a\rangle+\varepsilon^{\prime}_{t+1}$, where $\varepsilon^{\prime}_{t+1}$ is another realization of the noise. Statistically, a specific instance of $\langle\zeta_{t+1},a\rangle+\varepsilon_{t+1}$ and $\varepsilon^{\prime}_{t+1}$ are close to each other, when $\delta$ is small, albeit the resulting (small) mean-shift due to the drift in the governing parameter. For practical purposes, the impact of the drift can be absorbed into the noise term $\varepsilon_{t+1}$ when $\delta$ is small. As a result, one expects an algorithm tailored to handle piecewise stationarity to perform reasonably well for slowly drifting environments. On the other hand, if $\delta$ is large, the bias induced by $\zeta_{t+1}$ is large enough to disallow absorbing it into the zero-mean noise term. Over a few time-steps, the cumulative effect of this compounding bias is then large enough to completely violate the stationarity assumption. With large enough $\delta$, the change in $\theta_{t}$ over a few time-steps can be considered large enough to trigger a restart.
5 Summary and Outlook
In this work, we introduced DAL, a practical, prior-free black-box framework for NS bandits. Its plug-and-play design allows seamless integration with a wide range of stationary bandit algorithms. Through extensive experiments in both PS and drifting settings—spanning synthetic and real-world benchmarks—DAL consistently outperforms all prior-free baselines, including the black-box gold standard MASTER and the state-of-the-art ADA-OPKB. DAL even surpasses leading prior-based methods like WeightUCB and WGP-UCB. Finally, DAL achieves the strongest performance in real-world scenarios, which underscores its value as a practical and effective solution.
To explain these results, we provided insights on its theoretical nature. Specifically, using existing theoretical results and providing novel techniques, we showed that DAL achieves order-optimal regret in piecewise stationarity if its input stationary bandit algorithm has order-optimal stationary regret guarantees, under a mild assumption on the separatedness of the change-points. Crucially, DAL preserves and adapts the regret properties of its stationary input algorithm to the PS setting. As a result, it matches the best existing bounds in PS-LBs and PS-GLBs while improving the best known bounds for PS-SCBs and PS-KBs. Regarding drifting non-stationarity, we hypothesized key conditions under which DAL excels–an insight further validated through additional experiments under drifting settings. Our results suggest that a well-designed algorithm for the PS setting can extend to a broad range of drifting scenarios, bridging the gap between these two regimes.
While our work makes significant progress, it also opens several new directions. First, theoretical guarantees for change detection-based methods under drifting environments remain largely unexplored. Second, the current regret bounds for DAL rely on a separation condition between change-points—a standard assumption in the detection-based literature (see e.g., [5, 7, 27])—which nonetheless limits the extent to which DAL achieves fully prior-free theoretical optimality. Addressing these limitations could advance our understanding of detection-based methods in more continuous forms of non-stationarity. Finally, DAL’s modular nature invites future extensions to even richer settings, including general NS reinforcement learning. We believe that deepening the study of piecewise stationarity may be the key to tackling these broader challenges and DAL can serve as a solid foundation towards that goal.
Appendix A Experimental Details
A.1 Change Detectors
Below, we provide the algorithmic implementation of the Generalized Likelihood Ratio (GLR) test, including the necessary information for both its Gaussian and Bernoulli versions.
Algorithm 2 Generalized Likelihood Ratio (GLR) Test
Input: observations $X_{1},\dots,X_{n}$, false alarm probabililty $\delta_{\mathrm{F}}$, missed detection probability $\delta_{\mathrm{D}}$
Output: detection if a change-point is detected, else no-detection
1:  for $s=1$ to $n-1$ do
2:     Compute empirical means $\hat{\mu}_{1:s}$, $\hat{\mu}_{s+1:n}$, and $\hat{\mu}_{1:n}$.
3:     Compute test statistic:
$\mathrm{GLR}_{s}\leftarrow s\cdot\mathrm{kl}(\hat{\mu}_{1:s},\hat{\mu}_{1:n})+(n-s)\cdot\mathrm{kl}(\hat{\mu}_{s+1:n},\hat{\mu}_{1:n})$
4:     if $\mathrm{GLR}_{s}\geq 6\log\left(1+\log(n)\right)+\frac{5}{2}\log\left(\frac{4n^{3/2}}{\delta_{\mathrm{F}}}\right)+11.$ then
5:        return  detection
6:     end if
7:  end for
8:  return  no-detection
In the case of the Bernoulli GLR version we have that,
$\mathrm{kl}(x,y)=x\ln\left(\frac{x}{y}\right)+(1-x)\ln\left(\frac{1-x}{1-y}\right).$
The Bernoulli GLR test is used in the case of sub-Bernoulli distributions [7], e.g., distributions bounded in [0,1].
On the other hand, for the case of the Gaussian GLR version, assuming variance $\sigma^{2}$, we have that,
$\mathrm{kl}(x,y)=\frac{(x-y)^{2}}{2\sigma^{2}}.$
The Gaussian GLR test is employed in the case of $\sigma^{2}-$sub-Gaussian distributions, e.g., bounded distributions in $[a,b]$ for some $a<b\in\mathbb{R}$.
For the practical tuning of the detector, we follow the information provided in [27].
A.2 On Forced Exploration in Finite Action Spaces
In practice, the covering set $\mathcal{A}_{\mathrm{e}}$ would generally be selected according to Propositions 4.1 and 4.2, along with the specifications of Corollary 4.4. However, in finite-action settings, the full construction may not be feasible: $\mathcal{A}$ might not contain enough elements to satisfy the required conditions. For instance, in the PS-PB setting, $\mathcal{A}$ may not include $d$ linearly independent actions, while in the PS-KB case, it may lack the full $\delta_{T}$-covering net for the chosen $\delta_{T}$ in Corollary 4.4.
One might expect that when $|\mathcal{A}_{\mathrm{e}}|<d$ in PS-PBs or $|\mathcal{A}_{\mathrm{e}}|<\gamma_{T}$ in PS-KBs, the inability to detect all possible changes would degrade DAL’s performance. In general, however, the algorithm does not need to restart if the changes in the reward function does not change the mean reward obtained from each action. Crucially, as discussed in Appendix B.5, DAL retains order-optimality even in these constrained regimes.
Accordingly, when $|\mathcal{A}|<d$ or $|\mathcal{A}|<\gamma_{T}$, we simply set $\mathcal{A}_{\mathrm{e}}=\mathcal{A}$. In our experiments, the action set is finite (typically in the hundreds). For PS-PBs, the random generation of actions almost always guarantees $d$ linearly independent vectors. For PS-KBs, since $\gamma_{T}$ is typically large, we also use the full action set $\mathcal{A}$ as $\mathcal{A}_{\mathrm{e}}$ without impacting performance.
A.3 Real-World Data Preprocessing
Yahoo! R6A Dataset Construction.
For the NS bandit benchmark based on the Yahoo! R6A click log dataset [^5] Yahoo! Front Page Today Module User Click Log Dataset: https://webscope.sandbox.yahoo.com. we follow the main procedure provided in [10, 62]. We merge ten consecutive days of logs and we group the data by article ID and compute smoothed click-through rates (CTRs) using centered rolling averages over a 100-round window. This generates a time series of empirical CTRs for each article. We segment the dataset into ten distinct subperiods (each spanning half a day), filtering out actions with missing data or high noise. We further select a set of common actions present in all segments to ensure consistent tracking. We average CTRs within each subperiod and smoothing small deviations below a threshold $0.005$. We stack selected actions across multiple days into a single $K\times T$ matrix, where $K$ is the number of valid actions and $T$ the compressed time horizon. To reduce spurious noise and compress the time scale, we apply local smoothing. Finally, we apply post-processing filters to remove (i) globally high-value actions (outliers with inflated CTRs), and (ii) actions that persist as best for too many segments.
Figure 4: Mean rewards for the Yahoo! R6A dataset.
Stock Market Data Construction.
Regarding the stock market experiments we follow the procedure of [17]. For the first experiment, we use the data provided in [17]. For the other experiment, we collect daily closing prices of NASDAQ-100 companies using the Yahoo Finance API [^6] Data retrieved from Yahoo Finance using the publicly available yfinance package. Used solely for non-commercial, academic research purposes.. We filter out stocks with fewer than $T=2000$ trading days and align all time series over the most recent $T$ dates. From this pool, we remove stocks with extremely high volatility or mean price to make the problem non-trivial, then select the top $K$ most volatile stocks from the remainder. In both cases, the stock prices are scaled accordingly to lie in $[0,1]$. Each selected company’s scaled closing price series defines the mean-reward sequence for one arm in a $K$-armed bandit problem. Finally, we corrupt the reward at each time step with $\mathcal{N}(0,0.01)$ noise.
Figure 5: Daily closing prices from the dataset of [17].
Figure 6: Daily closing prices obtained from Yahoo Finance.
Live Traffic Data Construction.
We construct a NS bandit environment based on the Criteo live traffic dataset [18], following the preprocessing approach of [45] but modeling the problem as an NS-GLB rather than an NS-LB. Specifically, the dataset includes banners shown to users, associated contextual variables, and whether each banner was clicked. We retain the categorical variables cat1 through cat9, along with campaign, which uniquely identifies each campaign. These categorical features are one-hot encoded, and a dimensionality reduction via Singular Value Decomposition selects 50 resulting features. The parameter vector $\theta^{\star}$ is estimated using logistic regression. Rewards are then generated from this regression model with added Gaussian noise of variance $\sigma^{2}=0.01$. Unlike [45], in which the authors employ a single change, we introduce shifts in $\theta^{*}$ via a geometric change-point model with parameter $\xi=0.8$, chaging $60\%$ of the $\theta^{*}$ coordinates at each time-step to $-\theta^{*}$ and extend the horizon to $T=50000$.
Sensor Correlation Data Construction.
We use the Bioliq dataset from [31], comprising a week of readings from 20 power plant sensors. Following their setup, we construct an NS-SCB environment with 190 actions: the reward is $1$ if the last 1000 measurements exceed 2.04, and 00 otherwise. Evaluation is based on cumulative reward. Data available at https://github.com/edouardfouche/G-NS-MAB/tree/master/data.
Figure 7: Raw rewards obtained from the Bioliq dataset [31].
Ad Recommendation Data Construction.
We evaluate on the Zozo environment, a real-world ad recommender system from [46], using the preprocessed dataset of [31]. We construct an NS-GLB environment with all 80 ads (unlike their 10-action setup), assigning reward $1$ to any ad clicked within one second, and 00 otherwise. Evaluation is based on cumulative reward. Data available at https://github.com/edouardfouche/G-NS-MAB/tree/master/data.
Figure 8: Raw rewards obtained from the Zozo dataset [31].
A.4 Hardware Specifications
All experiments were employed on a desktop using an Intel(R) Xeon(R) W-2245 processor with 32 GB RAM. Each experiment had a total runtime below one hour.
Appendix B Theoretical Results
B.1 Regret Bounds of DAL in Piecewise Stationary Environments
As discussed in Section 4.2 of the paper, using Corollary 4.4, we can select different algorithms as input for DAL to attain or improve the state-of-the-art regret bounds in PS environments. Combining DAL with different bandit algorithms leads to the results of Table 1. It is evident that DAL matches the state-of-the-art regret bounds in PS-LBs and PS-GLBs, and DAL improves the best known bounds in the PS-SCB and PS-KB settings. Note that for PS-SCBs, the strongest result correspond to the prior-based WeightUCB [57]. As demonstrated in the final columns of the table, the order-wise dependence on problem parameters from the stationary setting seamlessly transfers to the PS setting without degradation.
Table 1: Regret bound comparison of algorithms for PS bandits, under the assumption of [27]. $\dagger$” denotes settings with finite number of actions, while MASTER, ADA-OPKB and SCB-WeightUCB also recover the appropriate bounds in this setting. $\bullet$” indicates prior-based algorithms.
PS Setting Non-Stationary Algorithm NS Algorithm Regret Bound in $\tilde{\mathcal{O}}(\cdot)$ DAL Input Regret Bound in $\tilde{\mathcal{O}}(\cdot)$
PS-LB MASTER [58] + LinUCB $d\sqrt{TN_{T}}$
ADA-OPKB [26] $d\sqrt{N_{T}T}$
DAL (ours) + LinUCB [1] $d\sqrt{N_{T}T}$ $d\sqrt{T}$
DAL (ours) + LinTS [2] $d^{3/2}\sqrt{N_{T}T}$ $d^{3/2}\sqrt{T}$
DAL (ours) + PEGE [33] $\sqrt{dN_{T}T}$ $\sqrt{dT}$
PS-GLB MASTER [58] + GLM-UCB $d\sqrt{N_{T}T}$
DAL (ours) + GLM-UCB [22] $d\sqrt{N_{T}T}$ $d\sqrt{T}$
DAL (ours) + GLM-TSL [32] $d^{3/2}\sqrt{N_{T}T}$ $d^{3/2}\sqrt{T}$
DAL (ours) + SupCB-GLM [37] $\sqrt{dN_{T}T}$ $\sqrt{dT}$
PS-SCB SCB-WeightUCB [57] $d^{2/3}T^{2/3}N_{T}^{1/3}$
DAL (ours) + OFU-ECOLog [20] $d\sqrt{N_{T}T}$ $d\sqrt{T}$
DAL (ours) + OFUL-MLogB [60] $d\sqrt{N_{T}T}$ $d\sqrt{T}$
DAL (ours) + OFUGLB [34] $d\sqrt{N_{T}T}$ $d\sqrt{T}$
PS-KB MASTER [58] + GPUCB $\gamma_{T}\sqrt{N_{T}T}$
ADA-OPKB [26] $\sqrt{d\gamma_{T}N_{T}T}$
DAL (ours) + GPUCB [14] $\gamma_{T}\sqrt{N_{T}T}$ $\gamma_{T}\sqrt{T}$
DAL (ours) + REDS [47] $\sqrt{\gamma_{T}N_{T}T}$ $\sqrt{\gamma_{T}T}$
B.2 Proof of Proposition 4.1
In the NS-PB setting, the reward at time $t$ is given by $f_{t}(a)=\mu(\langle\theta_{t},a\rangle)$ for all $a\in\mathcal{A}$, where $\mu$ is injective and $\theta_{t}\in\mathbb{R}^{d}$. To detect any changes in $\theta_{t}$, it suffices to detect changes in the values $\langle\theta_{t},a\rangle$ for a suitable set of actions.
Since $\mu$ is injective, each observation $y_{t,i}=\mu(\langle\theta_{t},a_{i}\rangle)$ can be inverted to recover the inner product:
$\langle\theta_{t},a_{i}\rangle=\mu^{-1}(y_{t,i}).$
Hence, observing $y_{t,i}$ is equivalent to observing $\langle\theta_{t},a_{i}\rangle$.
Suppose that $\mathcal{A}_{\mathrm{e}}\subseteq\mathcal{A}$ is the maximal linearly independent subset of $\mathcal{A}$. Then, the vector $\theta_{t}$ is uniquely determined by the inner products $\langle\theta_{t},a\rangle$ for $a\in\mathcal{A}_{\mathrm{e}}$. Therefore, any change in $\theta_{t}$ results in a detectable change in the vector of observations $(y_{t,i})_{a_{i}\in\mathcal{A}_{\mathrm{e}}}$, meaning that $\mathcal{A}_{\mathrm{e}}$ can be taken to be any maximal linearly independent subset of $\mathcal{A}$, with $\lvert\mathcal{A}_{e}\rvert\leq d$.
B.3 Proof of Proposition 4.2
In this subsection, we establish the construction of $\mathcal{A}_{e}$ in the NS-KB setting. According to Lemma 5 from [15], we have that every $f\in H_{k}$ with $\|f\|_{H_{k}}\leq B$ is Lipschitz continuous, satisfying the following,
$|f(x)-f(y)|\;\leq\;B\,L_{u}\,\lVert x-y\rVert_{2},\>\forall\,x,y\in\mathcal{A},\quad\text{where}\>L_{u}:=\sup_{z\in D}\max_{i,j\leq d}\Bigl{[}\frac{\partial^{2}k(p,q)}{\partial p_{i}\,\partial q_{j}}\Bigr{]}^{1/2}_{\;p=q=z}.$
Recall that $\mathcal{P}_{T}$ corresponds to the set of centers of the balls of an arbitrary $\delta_{T}$-cover of $\mathcal{A}\subseteq[0,R]^{d}$. Let $[a]_{\mathrm{e}}$ denote the action in $\mathcal{P}_{T}$ that is the closest to $a\in\mathcal{A}$, i.e., $[a]_{\mathrm{e}}=\mathrm{argmin}_{x\in\mathcal{P}_{T}}\|a-x\|_{2}$. Then, we can leverage the Lipschitz property of functions in the RKHS to obtain the following upper bound: For any $a\in\mathcal{A}$ and $f\in H_{k}$ with $\|f\|_{H_{k}}\leq B$,
$|f(a)-f([a]_{\mathrm{e}})|\overset{(a)}{\leq}BL_{u}\|a-[a]_{\mathrm{e}}\|_{2}\overset{(b)}{\leq}BL_{u}\delta_{T}.$ (1)
Step $(a)$ follows from the Lipschitz property in Lemma 5 of [15], and step $(b)$ results from the definition of a $\delta_{T}$-cover. Then, for any arbitrary functions $f$ and $f^{\prime}$ in $H_{k}$ with $\|f\|_{H_{k}},\|f^{\prime}\|_{H_{k}}\leq B$ and action $\tilde{a}\in\mathcal{A}$, we have
$|f([\tilde{a}]_{e})-f^{\prime}([\tilde{a}]_{e})|$ $\geq|f(\tilde{a})-f^{\prime}(\tilde{a})|-|f(\tilde{a})-f([\tilde{a}]_{\mathrm{e}})|-|f^{\prime}(\tilde{a})-f^{\prime}([\tilde{a}]_{\mathrm{e}})|$
$\overset{(a)}{\geq}|f(\tilde{a})-f^{\prime}(\tilde{a})|-2BL_{u}\delta_{T}$
>$\overset{(b)}{>}0$
where step $(a)$ is due to (1), and step $(b)$ is due to the assumption in Proposition 4.2. This indicates that the value of the reward function at $[\tilde{a}]_{\mathrm{e}}$ must change by a non-zero amount. Thus, one can use observations from action $[\tilde{a}]_{\mathrm{e}}$ in order to deduce whether the reward function has changed its value in action $\tilde{a}$. In addition, by the upper bound on the covering number, the cardinality of $\mathcal{P}_{T}$ is upper bounded by $\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$.
B.4 Sketch of Proof of Theorem 4.3
The proof of Theorem 4.3 follows exactly those of Theorem 1 and Corollary 1 in [27], with $A$ replaced by $N_{\mathrm{e}}$, due to the different number of actions in the covering set. For completeness, we provide a proof sketch of Theorem 4.3: First, we partition the regret into two cases. If no false alarm occurs and all changes are detected within a short delay, we can separate the regret into three components: the regret due to forced exploration, the regret during the short detection (restart) delay after changes, and the regret incurred by the stationary bandit algorithm after the change is detected. If not, we use a crude linear bound to bound the regret and show that the probability of false alarm and that of late detection are low, which ensures that the regret due to detection failure is small.
B.5 Proof of Corollary 4.4
In PS-PBs, $N_{\mathrm{e}}=d$, $p\geq 1/2$, and $q=0$. Thus, $R_{T}=\tilde{\mathcal{O}}(\sqrt{dN_{T}T}+d^{p}\sqrt{N_{T}T})=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T})$.
In PS-KBs, $p,q\geq 1/2$. We can upper bound $N_{\mathrm{e}}$ using the fact that $|\mathcal{P}_{T}|\leq\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$. Thus, $N_{e}\leq\lceil C\gamma_{T}^{2q/d}\rceil^{d}$ with $\delta_{T}=\frac{Rd^{1/2-2p/d}}{2(C\gamma_{T}^{2q})^{1/d}}$ and $R_{T}=\tilde{\mathcal{O}}(\sqrt{d^{2p}\gamma_{T}^{2q}N_{T}T}+d^{p}\gamma_{T}^{q}\sqrt{N_{T}T})=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T})$.
We emphasize that when the number of action is smaller than the covering number, i.e., $|\mathcal{A}|<\lceil C\gamma_{T}^{2q/d}\rceil^{d}\leq\gamma_{T}$, then we can set $\mathcal{A}_{e}$ to be the entire action set $\mathcal{A}$. In this case, $N_{\mathrm{e}}<\gamma_{T}$, guaranteeing order-optimal regret.